OpenStudy (anonymous):

(2)/(3)-(1)/(3x)+(7)/(5x)

thomaster (thomaster):

why is everything in parentheses?

thomaster (thomaster):

Is it like this? \[\frac{ 3 }{ 2}-\frac{ 1 }{ 3x}+\frac{ 7 }{ 5x }\]or\[\frac{ \frac{ 2 }{ 3}-1 }{ 3x+\frac{ 7 }{ 5x } }\]or\[\frac{ \frac{ 2 }{ 3-1 } }{ \frac{ 3x+7 }{ 5x } }\]